# B. Nikolić

### v1

@@ Abstract
In this article we consider type II superstring in the pure spinor formulation with constant background fields in the context of T-dualization. First we prove that bosonic and fermionic T-dualization commute using already known T-dual transformation laws for bosonic and fermionic T-dualization. Consequently, the T-dual transformation laws of the full T-dualization are obtained. At the end the full T-dualization is realized in double space and it is showed that Buscher procedure and double space approach are equivalent in this specific case.
@@ 1 Introduction
Our focus in this article will be on T-duality transformations. The simplest and most fundamental picture of T-duality transformation is a compactification of one spatial dimension on a circle of radius $R$. For closed string along compactified direction $X$ the following periodic condition must hold, $X(\sigma+2\pi)=X(\sigma)+2\pi RW$, where $W$ is the winding number. This periodic condition implies that momentum along that direction is quantizied, $P=K/R$, where $K$ is an integer named Kaluza-Klein excitation number. It turns out that theory, after compactification, becomes invariant under exchanges, $W\leftrightarrow K$ and $R\leftrightarrow\alpha^{\prime}/R$, where $\alpha^{\prime}$ is Regge slope parameter. In this way we show the physical equivalence of the theories on the large and small scales, which is the basic notion of T-duality.
@@
Analytical form of the T-duality transformation is given by Buscher procedure [2], which is considered in literature as a definition of T-duality. The starting point of this approach is identifying the direction $X$ along which the global shift symmetry exists, $X\rightarrow X+a$. Further we localize the observed symmetry introducing covariant world-sheet derivative, $\partial_{\pm}X\rightarrow D_{\pm}X=\partial_{\pm}X+v_{\pm}$, where $v_{\pm}$ is a gauge field. Then we fix the gauge condition putting $X=const.$. In order for the initial and transformed (T-dualized) theories to have the same number of degrees of freedom, one has to add a term with a Lagrange multiplier $y$ to the initial action. On the equation of motion for $y$ the initial theory is restored, while on the equations of motion for $v_{\pm}$ we obtain T-dual theory. One of the important outcomes of the procedure is analytical form of the relation between initial coordinates and T-dual coordinates - T-dual transformation laws. Buscher procedure is applicable on the models with constant backgrounds as well as on the cases where background fields do not depend on some coordinates [3, 4, 5, 6, 7, 8].
@@
The superstring theories have both bosonic and fermionic coordinates [1]. Formally, Buscher procedure can be generalized for the fermionic coordinates and that is so called fermionic T-duality [9]. Unlike bosonic T-duality, fermionic T-duality involves shifting along fermionic isometry directions. It is a non-local and formal operation, only defined at the level of the classical string sigma model and requiring a careful treatment of global and quantum consistency conditions.
@@
Model used in this paper is type II superstring in pure spinor formulation [10]. In general, the model is pretty complex and all background fields and their field strengths depend on all coordinates $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$, where $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are fermionic coordinates. Because of that we make a strong approximation, in accordance with the consistency conditions, taking all background fields constant as well as Ramond-Ramond field strength.
@@
The starting point are the papers [11] where we analyzed separately in details bosonic and fermionic T-duality using pure spinor formulation of the type II superstring theory with constant background fields. We obtained the form of T-dual theory as well as T-dual transformation laws, which relates initial and T-dual coordinates. Also we represented both bosonic and fermionic T-duality in the double space [12], space spanned by initial and T-dual coordinates (Lagrange multipliers). In double space T-dualization is realized through an appropriate permutation matrix and it acts as symmetry transformation. We proved the equivalence of the Buscher analytical approach and algebraic double space approach for this specific choice of action.
@@
Also the double space representation of the full T-dualization was done. The double space coordinate is of the form $Z^{A}=(x^{\mu},\theta^{\alpha},\bar{\theta}^{\alpha},y_{\mu},z_{\alpha},\bar{z}_{\alpha})^{T}$, where $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are the initial coordinates and $y_{\mu}$, $z_{\alpha}$ and $\bar{z}_{\alpha}$ T-dual coordinates. Permutation matrix $\mathcal{T}^{A}{}_{B}$ is chosen in such a form to exchange the set of initial coordinates with the set of the T-dual coordinates and produces T-dual double coordinate, $^{\star}Z^{A}=\mathcal{T}^{A}{}_{B}Z^{B}$. Form of the action induces that there two generalized metrics, one for each chiral sector, $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$. Demanding that T-dual double coordinate has the same T-dual transformation law as initial one, we get the form of the T-dual background fields and confirm the agreement with the analytical Buscher approach.
@@ 2 Choice of action and brief review of the bosonic and fermionic T-duality
Let us explain the notation and conventions. Here the vectors $X^{M}$ and $X^{N}$ are given as
@@
Implementation of the combined bosonic and fermionic T-duality is possible if there exists at least two isometry directions - one in the bosonic and one in fermionic sector. Consequently, one has to impose some restrictions. Trivial solution for bosonic T-dualization is to eliminate the explicit dependence on $x^{\mu}$ and keep only dependence on $\partial_{\pm}x^{\mu}$. Some specific models can contain $x^{\mu}$ dependence (see [14]) where isometry still exists. But here I chose the trivial solution which is in accordance withe the consistency conditions from Ref.[13]. For fermionic T-duality the situation is pretty similar. One has to avoid dependence on $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$. Such assumption produces the following consequence - the auxiliary superfields must be zero which is clear from the consistency conditions given in Ref.[13]. Choosing physical background fields is not problematic in the sense of the consistency conditions [13] except Ramon-Ramond (RR) field strength ${\rm P}^{\alpha\beta}$. There is a possibility of back reaction to the rest background fields. But detailed analysis presented in Ref.[15] (subsection 2.1) showed that choice of constant RR field strength is correct.
@@ 2.3 A brief review of the bosonic T-duality
The world-sheet derivatives $\partial_{\pm}x^{\mu}$ are replaced with the appropriate covariant ones $D_{\pm}x^{\mu}=\partial_{\pm}x^{\mu}+v^{\mu}_{\pm}$, where $v_{\pm}^{\mu}$ are the gauge fields. Fixing the gauge, $x^{\mu}=const.$, the covariant derivative drives into gauge field. In order to make $v^{\mu}_{\pm}$ to be unphysical degrees of freedom, we add to the action the term
@@
where $y_{\mu}$ are Lagrange multipliers (T-dual coordinates). Finally, the gauge fixed action is getting the form
@@ 3.1 First bosonic then fermionic T-dualization
Let us consider the case where the bosonic T-dualization is done first and then the fermionic one. It is sufficient to consider only one subset of the T-dual transformation laws, for example, the relations where derivatives of the initial coordinates are expressed in terms of the derivatives of the undulized coordinates.
@@ 3.2 First fermionic then bosonic T-dualization
As we see the full T-dual transformation law for $\partial_{+}x^{\mu}$ given in Eqs.(3.7) and (3.9) are the same, which suggests that bosonic and fermionic T-dualization in this specific case commute. For full confirmation it remains to check the rest of T-dual transformation laws.
@@
Taking into account the expressions for bosonic T-dualized gravitino fields (2.46), we obtained the expressions same as in Eq.(3.3). The T-dual transformation laws after full T-dualization for $\partial_{-}\bar{\theta}^{\alpha}$ and $\partial_{+}\theta^{\alpha}$ are given in Eqs.(2.49) and (2.51), because additional BRST gauge term is not affected by the bosonic T-dualization.
Finally, we proved that bosonic and fermionic Buscher T-dualization procedures in this specific choice of background fields commute. Let us note that commutation can be proved just on the level of the dualized background fields. But, in this way, we proved the commutation and, simultanously, obtained the full T-dual transformation laws. Further it is easy to find the expressions for full T-dualized background fields
@@
where $\star$ in the supscript denotes full T-dualization.
@@ 4 Double space representation of the full T-dualization
In the previous section we found the expressions for T-dual transformation laws in which the world-sheet derivatives of the initial coordinates $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are related with the derivatives of the coordinates of the full T-dualized theory $y_{\mu}$, $z_{\alpha}$ and $\bar{z}_{\alpha}$. Here we want to rewrite the analytical T-dualization process in the double space. In order to accomplish that goal, first we have to find expressions for the inverse transformations.
@@ 4.1 Inverse T-dual transformation laws for full T-dualization
Let us start with the transformations for $\partial_{+}x^{\mu}$ (3.9) and $\partial_{+}\bar{\theta}^{\alpha}$ (3.12) and multiply the first one with $\bar{\Psi}^{\alpha}{}_{\mu}$. Adding obtained expressions and using the definition of $\tilde{P}^{\alpha\beta}$ (2.28), the following transformation is obtained
@@
Inserting these expressions into expressions for the full T-dual transformation law (3.9) and multiplying by $\Pi_{+\mu\rho}$, we get the inverse relation
@@
Applying the same procedure on the opposite chiral sector, we have the other relation
@@
The rest two transformations fullfilling this set of transformations are given in Eqs.(2.49) and (2.51).
@@ 4.2 Double space representation of the full T-dualization
Here we will rewrite the full T-dualization (bosonic+fermionic) in terms of the double space which is spanned by
@@
From now on when we say double space coordinate we think on $Z^{A}$ given in (4.8). The full T-dualized double coordinate is defined as
@@
First we have to rewrite the T-dual transformation laws for full T-dualization in the appropriate form. Let us start with the T-dual law for $\partial_{+}x^{\mu}$ (3.9) and multiply from the right with
@@
As a result, the folllowing equation is obtained
@@
where we used the definition of $\Theta^{\mu\nu}_{-}$ (2.30). Putting $\Pi_{-\mu\nu}=B_{\mu\nu}-\frac{1}{2}G_{\mu\nu}$ in the first term on the right-hand side of the equation, after short calculation, the following equation is obtained
@@
In the inverse relation (4.4), using the same logic, we insert
@@
To complete the set of equations, we rewrite the already obtained equations in the slaightly changed form
@@
Consequently, T-dual generalied metrics, $\mathcal{H}_{\pm MN}$, transform as
@@
From this equation, it is possible to find the background fields of the full T-dualized theory. Our goal is to compare them with those obtained using analytical method and confirm applicability of double space formalism in this case.
Let us analyze ++ chiral sector. Taking into account that matrix (4.26) is big one, let us simplify it as
@@
From the equality of block compenents $^{\star}B=C$, we have
@@
Using the simple fact that $\frac{1}{2}{}^{\star}G{}^{\star}G^{-1}=\frac{1}{2}\check{G}^{-1}\check{G}$, we get the following useful relation
@@
Here we used the fact that $\check{\Theta}_{+}=-\check{\Theta}_{-}^{T}$.
Let us go back to the equation (4.34). The T-dual effective metric $^{\star}G_{E}^{\mu\nu}$ is defined as
@@
Now using the auxiliary relation (4.39), we have
@@
At the end, using (4.34), we are coming finally to the result
@@
As the undetermined T-dual background field remained $^{\star}B^{\mu\nu}$. Using expression for $^{\star}G^{\mu\nu}$ and relation (4.38), we get
@@
Comparing T-dual background fields obtained in the analytical approach (3.13)-(3.15) with the expressions above obtained in the double space approach, we see the full accordance. The same expressions would be obtained analyzing $\mathcal{H}_{-MN}$.

### v2

@@ Abstract
In this article I consider type II superstring in the pure spinor formulation with constant background fields in the context of T-dualization. First I prove that bosonic and fermionic T-dualization commute using already known T-dual transformation laws for bosonic and fermionic T-dualization. Consequently, the T-dual transformation laws of the full T-dualization are obtained. At the end the full T-dualization is realized in double space and it is showed that Buscher procedure and double space approach are equivalent in this specific case.
@@ 1 Introduction
My focus in this article will be on T-duality transformations. The simplest and most fundamental picture of T-duality transformation is a compactification of one spatial dimension on a circle of radius $R$. For closed string along compactified direction $X$ the following periodic condition must hold, $X(\sigma+2\pi)=X(\sigma)+2\pi RW$, where $W$ is the winding number. This periodic condition implies that momentum along that direction is quantizied, $P=K/R$, where $K$ is an integer named Kaluza-Klein excitation number. It turns out that theory, after compactification, becomes invariant under exchanges, $W\leftrightarrow K$ and $R\leftrightarrow\alpha^{\prime}/R$, where $\alpha^{\prime}$ is Regge slope parameter. In this way I show the physical equivalence of the theories on the large and small scales, which is the basic notion of T-duality.
@@
Analytical form of the T-duality transformation is given by Buscher procedure [2], which is considered in literature as a definition of T-duality. The starting point of this approach is identifying the direction $X$ along which the global shift symmetry exists, $X\rightarrow X+a$. Further the observed symmetry is localized introducing covariant world-sheet derivative, $\partial_{\pm}X\rightarrow D_{\pm}X=\partial_{\pm}X+v_{\pm}$, where $v_{\pm}$ is a gauge field. Then the gauge condition is fixed putting $X=const.$. In order for the initial and transformed (T-dualized) theories to have the same number of degrees of freedom, one has to add a term with a Lagrange multiplier $y$ to the initial action. On the equation of motion for $y$ the initial theory is restored, while on the equations of motion for $v_{\pm}$ T-dual theory is obtained. One of the important outcomes of the procedure is analytical form of the relation between initial coordinates and T-dual coordinates - T-dual transformation laws. Buscher procedure is applicable on the models with constant backgrounds as well as on the cases where background fields do not depend on some coordinates [3, 4, 5, 6, 7, 8].
@@
The superstring theories have both bosonic and fermionic coordinates [1]. Formally, Buscher procedure can be generalized for the fermionic coordinates and that is so called fermionic T-duality [9]. In the first paper of [9], Berkovits and Maldacena showed that tree level superstring theories in presence of supersymmetric background fields posses a new kind of symmetry. They analyzed the gluon scattering amplitudes in N = 4 super Yang-Mills theory and discovered that new symmetry - fermionic T-duality. It is a part of the dual superconformal symmetry which should be connected to integrability and it is valid just at string tree level. Unlike bosonic T-duality, fermionic T-duality involves shifting along fermionic isometry directions. It is a non-local and formal operation, only defined at the level of the classical string sigma model and requiring a careful treatment of global and quantum consistency conditions.
@@
Model used in this paper is type II superstring in pure spinor formulation [10]. In general, the model is pretty complex and all background fields and their field strengths depend on all coordinates $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$, where $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are fermionic coordinates. Because of that I make a strong approximation, in accordance with the consistency conditions, taking all background fields constant as well as Ramond-Ramond field strength.
@@
The starting point are the papers [11] where bosonic and fermionic T-duality are analyzed separately in details using pure spinor formulation of the type II superstring theory with constant background fields. T-dual theory is obtained as well as T-dual transformation laws, which relates initial and T-dual coordinates. Also both bosonic and fermionic T-duality are represented in the double space [12], space spanned by initial and T-dual coordinates (Lagrange multipliers). In double space T-dualization is realized through an appropriate permutation matrix and it acts as symmetry transformation. The equivalence of the Buscher analytical approach and algebraic double space approach for this specific choice of action is proved.
@@
Also the double space representation of the full T-dualization was done. The double space coordinate is of the form $Z^{A}=(x^{\mu},\theta^{\alpha},\bar{\theta}^{\alpha},y_{\mu},z_{\alpha},\bar{z}_{\alpha})^{T}$, where $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are the initial coordinates and $y_{\mu}$, $z_{\alpha}$ and $\bar{z}_{\alpha}$ T-dual coordinates. Permutation matrix $\mathcal{T}^{A}{}_{B}$ is chosen in such a form to exchange the set of initial coordinates with the set of the T-dual coordinates and produces T-dual double coordinate, $^{\star}Z^{A}=\mathcal{T}^{A}{}_{B}Z^{B}$. Form of the action induces that there two generalized metrics, one for each chiral sector, $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$. Demanding that T-dual double coordinate has the same T-dual transformation law as initial one, I get the form of the T-dual background fields and confirm the agreement with the analytical Buscher approach.
@@ 2 Choice of action and brief review of the bosonic and fermionic T-duality
Let me explain the notation and conventions. Here the vectors $X^{M}$ and $X^{N}$ are given as
@@
Implementation of the combined bosonic and fermionic T-duality is possible if there exists at least two isometry directions - one in the bosonic and one in fermionic sector. Consequently, one has to impose some restrictions. Trivial solution for bosonic T-dualization is to eliminate the explicit dependence on $x^{\mu}$ and keep only dependence on $\partial_{\pm}x^{\mu}$. Some specific models can contain $x^{\mu}$ dependence (see [14]) where isometry still exists. But here I chose the trivial solution which is in accordance withe the consistency conditions from Ref.[13]. For fermionic T-duality the situation is pretty similar. One has to avoid dependence on $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$. Such assumption produces the following consequence - the auxiliary superfields must be zero which is clear from the consistency conditions given in Ref.[13]. Choosing physical background fields is not problematic in the sense of the consistency conditions [13] except Ramond-Ramond (RR) field strength ${\rm P}^{\alpha\beta}$. There is a possibility of back reaction to the rest background fields. But detailed analysis presented in Ref.[15] (subsection 2.1) showed that choice of constant RR field strength is correct.
@@ 2.3 A brief review of the bosonic T-duality
The world-sheet derivatives $\partial_{\pm}x^{\mu}$ are replaced with the appropriate covariant ones $D_{\pm}x^{\mu}=\partial_{\pm}x^{\mu}+v^{\mu}_{\pm}$, where $v_{\pm}^{\mu}$ are the gauge fields. Fixing the gauge, $x^{\mu}=const.$, the covariant derivative drives into gauge field. In order to make $v^{\mu}_{\pm}$ to be unphysical degrees of freedom, the term
@@
is added to the action, where $y_{\mu}$ are Lagrange multipliers (T-dual coordinates). Finally, the gauge fixed action is getting the form
@@ 3.1 First bosonic then fermionic T-dualization
Let me consider the case where the bosonic T-dualization is done first and then the fermionic one. It is sufficient to consider only one subset of the T-dual transformation laws, for example, the relations where derivatives of the initial coordinates are expressed in terms of the derivatives of the undualized coordinates.
@@ 3.2 First fermionic then bosonic T-dualization
It is obvious the full T-dual transformation law for $\partial_{+}x^{\mu}$ given in Eqs.(3.7) and (3.9) are the same, which suggests that bosonic and fermionic T-dualization in this specific case commute. For full confirmation it remains to check the rest of T-dual transformation laws.
@@
Taking into account the expressions for bosonic T-dualized gravitino fields (2.46), the expressions same as in Eq.(3.3) is obtained. The T-dual transformation laws after full T-dualization for $\partial_{-}\bar{\theta}^{\alpha}$ and $\partial_{+}\theta^{\alpha}$ are given in Eqs.(2.49) and (2.51), because additional BRST gauge term is not affected by the bosonic T-dualization.
Finally, it is proved that bosonic and fermionic Buscher T-dualization procedures for this specific choice of background fields commute. Notice that commutation can be proved just on the level of the dualized background fields. But, in this way, I proved the commutation and, simultaneously, obtained the full T-dual transformation laws. Further it is easy to find the expressions for full T-dualized background fields
@@
where $\star$ in the left superscript denotes full T-dualization.
@@ 4 Double space representation of the full T-dualization
In the previous section I found the expressions for T-dual transformation laws in which the world-sheet derivatives of the initial coordinates $x^{\mu}$, $\theta^{\alpha}$ and $\bar{\theta}^{\alpha}$ are related with the derivatives of the coordinates of the full T-dualized theory $y_{\mu}$, $z_{\alpha}$ and $\bar{z}_{\alpha}$. Here I want to rewrite the analytical T-dualization process in the double space. In order to accomplish that goal, first I have to find expressions for the inverse transformations.
@@ 4.1 Inverse T-dual transformation laws for full T-dualization
Let me start with the transformations for $\partial_{+}x^{\mu}$ (3.9) and $\partial_{+}\bar{\theta}^{\alpha}$ (3.12) and multiply the first one with $\bar{\Psi}^{\alpha}{}_{\mu}$. Adding obtained expressions and using the definition of $\tilde{P}^{\alpha\beta}$ (2.28), the following transformation is obtained
@@
Inserting these expressions into expressions for the full T-dual transformation law (3.9) and multiplying by $\Pi_{+\mu\rho}$, the inverse relation is obtained
@@
Applying the same procedure on the opposite chiral sector, I get the other relation
@@
The rest two transformations fulfilling this set of transformations are given in Eqs.(2.49) and (2.51).
@@ 4.2 Double space representation of the full T-dualization
Let me rewrite the full T-dualization (bosonic+fermionic) in terms of the double space which is spanned by
@@
From now on when I say double space coordinate I think on $Z^{A}$ given in (4.8). The full T-dualized double coordinate is defined as
@@
First I have to rewrite the T-dual transformation laws for full T-dualization in the appropriate form. Let me start with the T-dual law for $\partial_{+}x^{\mu}$ (3.9) and multiply from the right with
@@
As a result, the following equation is obtained
@@
where the definition of $\Theta^{\mu\nu}_{-}$ (2.30) is used. Putting $\Pi_{-\mu\nu}=B_{\mu\nu}-\frac{1}{2}G_{\mu\nu}$ in the first term on the right-hand side of the equation, after short calculation, the following equation is obtained
@@
In the inverse relation (4.4), using the same logic, one inserts
@@
To complete the set of equations, I rewrite the already obtained equations in the slightly changed form
@@
Consequently, T-dual generalised metric, $\mathcal{H}_{\pm MN}$, transform as
@@
From this equation, it is possible to find the background fields of the full T-dualized theory. The goal is to compare them with those obtained using analytical method and confirm applicability of double space formalism in this case.
Let the ++ chiral sector be the first for analysis. Taking into account that matrix (4.26) is big one, it is simplified as
@@
From the equality of block components $^{\star}B=C$, it follows
@@
Using the simple fact that $\frac{1}{2}{}^{\star}G{}^{\star}G^{-1}=\frac{1}{2}\check{G}^{-1}\check{G}$, I get the following useful relation
@@
Here the fact $\check{\Theta}_{+}=-\check{\Theta}_{-}^{T}$ is used.
Let me go back to the equation (4.34). The T-dual effective metric $^{\star}G_{E}^{\mu\nu}$ is defined as
@@
Now using the auxiliary relation (4.39), one gets
@@
At the end, using (4.34), finally one obtains the result
@@
As the undetermined T-dual background field remained $^{\star}B^{\mu\nu}$. Using expression for $^{\star}G^{\mu\nu}$ and relation (4.38), I get
@@
Comparing T-dual background fields obtained in the analytical approach (3.13)-(3.15) with the expressions above obtained in the double space approach, the full accordance is obvious. The same expressions would be obtained analyzing $\mathcal{H}_{-MN}$.